The question is multiply:
(x^8 + 7x) (x^8 - 7x)
Here's what I got.
x^8(x^8) + x^8(-7x) + 7x(x^8) + 7x(-7x)=
(x^16 - 49x^2
Is this right?

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

(x^8 + 7x) (x^8 - 7x)
Here's what I got.
x^8(x^8) + x^8(-7x) + 7x(x^8) + 7x(-7x)=
(x^16 - 49x^2
Is this right?
YES! Good work.

Answer by Alan3354(69443)

The question is multiply:
(x^8 + 7x) (x^8 - 7x)
Here's what I got.
x^8(x^8) + x^8(-7x) + 7x(x^8) + 7x(-7x)=
(x^16 - 49x^2
It's right. Add a right side parentheses, tho.